OpenStudy (anonymous):

modulus amplitude form of 1+sinx+icosx?

hartnn (hartnn):

amplitude of a+ib is \(\sqrt{a^2+b^2}\) so, ampl of 1+sinx+icosx will be \(\sqrt{(1+\sin x)^2+\cos^2 x}\) simplify,....